In 2000 Jengka Inc. issued bonds with an 8 percent coupon rate and a RM1,000 face
value. The bonds will mature on March 1, 2025. If an investor purchased one of
these bonds on March 1, 2012, determine the yield to maturity if the investor paid
RM1,100 for the bond.
1
Expert's answer
2015-03-25T10:02:51-0400
N = 25, if = 8%, F = RM1,000, on March 1, 2012 P = RM1,100.
The bond's current value is:
P = F*if((1 - (1 + i)^-n)/i) + F(1 + i)^-n
where:
C = F * iF = coupon payment
N = number of payments
i = market interest rate, or required yield, M = face value
P = market price of bond.
1100 = 80*((1 - 1/(1 + i)^25)/i) + 1000/(1 + i)^25
The yield to maturity i = 7.13%.